OpenStudy (anonymous):

Can someone explain how to solve this please! Medal! (2x – 1)(x + 7) = 0

OpenStudy (anonymous):

http://wolframalpha.com/

OpenStudy (ybarrap):

$$ (2x – 1)(x + 7) = 0\\ 2x\times x+2x\times 7-1\times x-1\times 7=0 $$ Does this make sense? Can you simplify?

OpenStudy (pawanyadav):

Put the two brackets separately =0 and solve

OpenStudy (ybarrap):

A simpler way is this $$ x+7=0\\ 2x-1=0 $$ Because in the product, if EITHER of these is zero, the whole thing is zero. So you will have TWO solutions. That means two different values for x.

OpenStudy (pawanyadav):

You will get two valves of x

OpenStudy (ybarrap):

For x+7=0 What value of x will make this equal to zero?

OpenStudy (pawanyadav):

Just substract 7 from both side

OpenStudy (pawanyadav):

Another value of x is 1/2 tell me how?

OpenStudy (pawanyadav):

Do you understand ?

OpenStudy (anonymous):

yes but can you do a summary of it all in one post im not really seeing it together

OpenStudy (pawanyadav):

First x+7=0 (substract 7 from both side) x=-7 Second 2x-1=0 (add 1 both side) 2x=1. (divide by 2 both sides) x=1/2 So you get x=-7 and1/2